Adam rode his skateboard 4/12 of a mile on Monday, 3/12 of a mile on Tuesday and 3/12 of a mile on Thursday. How far did he skate on those 3 days?
Answer by stanbon(75887)

Adam rode his skateboard 4/12 of a mile on Monday, 3/12 of a mile on Tuesday and 3/12 of a mile on Thursday. How far did he skate on those 3 days?
-----
(4+3+3)/12 = 10/12 = 5/6 miles
--------------
Cheers,
Stan H.
-----------